22. Find the critical points of the function and use the Second Derivative Test to determine whether each critical point corrcalc 3/multivariable calculus problem
Answer #1

((и, 2) = + 2ир – 25 -Ion би* 1 2 y - To fмг 6 fuy = 2 fox 24 - Чу f=2. £ 2-Ч fuso - 2 Зи“+ 2 - гово 2) Su“+ и - to = 0 23и“-